# Cubic fourfolds with birational Fano varieties of lines

### v1

@@ 1. Introduction
    Fourier-Mukai partnership, i.e. equivalence of the Kuznetsov components of the derived categories of the two cubics.
Motivated by the famous rationality conjectures for cubic fourfolds (see [hassett], [Kuznet], [addington], [MR4292740]), Huybrechts conjectured that (FM) implies (BE) [Huybrechtscubicsbook, Ch. 6, Conjecture 3.21]. A very general cubic fourfold, i.e. $X\not\in\cup\,\mathcal{C}_{d}$, has no non-trivial Fourier-Mukai partners by [Huy17]. This conjecture becomes interesting for $X\in\mathcal{C}_{d},$ with some evidence provided by Fan and Lai [FL24] who verified the conjecture for a very general cubic in $\mathcal{C}_{20}$. We provide further evidence for this conjecture by considering two special families of cubics in $\mathcal{C}_{12}$ (see Theorems 3.1 and 5.1). The first is the family $\mathcal{M}_{nonsyz}$, parametrizing cubics containing a non-syzygetic pair of cubic scrolls, first studied in [BFMQ24]. The second is a component of $\mathcal{C}_{8}\cap\mathcal{C}_{12}$, first studied in [BolognesiRusso] and [yang2021lattice] and more recently in [C8C12]. On the other hand, it is well known that (BE) does not imply (FM), for example by considering two very general cubics in $\mathcal{C}_{14}$.
@@
It has been conjectured that (FM) implies (DF) [Huybrechtscubicsbook, Ch. 6, Remark 3.28(iii)], for which we provide corroboratory evidence (see Theorems 3.1, 4.1, and 5.1). On the other hand, we prove that (DF) implies neither (FM) nor (BF), by considering the very general member of $\mathcal{C}_{32}$ (See Theorem 6.1), and $\mathcal{C}_{546}$ (see Theorem 7.1), respectively.
@@ 2.1. Special cubic fourfolds
have an associated K3 surface $S$, meaning $\widetilde{H}(X,\mathbb{Z})\cong\widetilde{H}(S,\mathbb{Z})$. All cubic fourfolds known to be rational belong to $\mathcal{C}_{d}$ for some $d$ satisfing ()(**)( ∗ ∗ ), leading to the famous conjecture that cubic fourfolds are rational if and only if they admit a labelling of disrciminant $d$ satisfying ()(**)( ∗ ∗ ). Kuznetsov conjectured in [kuzcubic] that $X$ is rational if and only if $\mathcal{A}_{X}\simeq D^{b}(S)$ for some K3 surface $S$, which was shown in [AddingtonTwoRatConjs] and [MR4292740] to coincide with the previous conjecture.
@@ 2.3. Fano varieties of lines
The variety $F$ is naturally equipped with the Plücker polarization $g\in\mathrm{NS}(F)$, with $g^{2}=6$ and divisibility $2$. Recall that $\mathcal{M}_{6}^{(2)}$ is the moduli space of hyperkähler fourfolds of K3[2] type equipped with a polarization $L$ with $L^{2}=6$ and divisibility $2$. There is a period map: $p\colon\mathcal{M}_{6}^{(2)}\rightarrow\mathcal{D}_{6}^{(2)},$ where $\mathcal{D}_{6}^{(2)}$ is the relevant period domain (see for instance [DebarreSurvey, Section 3.9]).
@@ 4.1. Lattice theory and the movable cone
Conversely, if $\varphi$ preserves the positive cone and has the prescribed action on $H,$ then one can extend $\varphi$ to an isometry $\tilde{\varphi}\in\mathrm{O}(H^{2}(F,\mathbb{Z}))$. (Here we are using the fact that $q$ does not represent $-2$, so the positive cone and movable cone coincide.) By construction it preserves the Hodge structure, and hence $\varphi\in\mathrm{Im}(\pi\colon\mathrm{Mon}^{2}_{Hdg}(F)\rightarrow\mathrm{O}(\mathrm{NS}(F))).$ By Theorem 2.1, the conclusion follows. ∎
@@
As previously mentioned, $q$ does not represent $-2$, so $\mathrm{Mov}(F)=\overline{\mathrm{Pos}(F)}$. Let $\rho=ag+b\lambda$ where $6a^{2}+4ab-6b^{2}$, so $\rho^{\perp}$ is a wall in $\mathrm{Mov}(F)$. If $|a|>1$, then $|13a+18b|<|a|$ or $|25a-18b|<|a|$, so exactly one of the classes $AB\cdot\rho$ or $(AB)^{-1}\cdot\rho$ has smaller first coordinate than $\rho$. Inductively, we see that $\rho=\pm(AB)^{n}\cdot(g+2\lambda)$ for a unique integer $n$. Hence $\mathrm{Bir}(F)$ acts freely and transitively on the walls of $\mathrm{Mov}(F)$. Further, $\mathrm{Bir}(F)$ acts transitively on the chambers of $\mathrm{Mov}(F)$ since $\mathrm{Mon}^{2}_{Hdg}\cong W_{Bir}\cong\mathrm{Bir}(F)$.
@@ 5.2. Geometry
  • since $[\ell]\not\in P^{*}$, $C_{\ell}\not\subset\Pi^{\prime}$, so the total transform of $C_{\ell}$ under the projection $Z_{T_{L}}\dashrightarrow X^{\prime}$ is a conic $C_{\ell}^{\prime}$.
Let $\ell^{\prime}$ be the line residual to $C_{\ell}^{\prime}$ in $X^{\prime}$. One easily checks that the formula $\pi([\ell])=[\ell^{\prime}]$ agrees with the composition of birational equivalences $F\dashrightarrow\widetilde{W}_{T_{L}}^{\perp}\dashrightarrow F^{\prime}$. So, $\pi\colon F\dashrightarrow F^{\prime}$ is regular away from $\mathbb{P}(V^{\vee})\cup P^{*}$. To see that $\pi$ is indeterminate along $\mathbb{P}(V^{\vee})\cup P^{*}$, it suffices to check that the birational maps $f\colon\widetilde{W}_{T_{L}}^{\perp}\dashrightarrow F$ and $f^{\prime}\colon\widetilde{W}_{T_{L}}^{\perp}\dashrightarrow F^{\prime}$ have different base loci. Following [BFMQ24, Lemma 3.5 and Proposition 3.6], a point $w\in\widetilde{W}_{T_{L}}^{\perp}$ parametrizes a singular quadric threefold $Q_{w}$, and $f$ (respectively $f^{\prime}$) is defined at $w$ if and only if $Q_{w}\cap\Pi=\emptyset$ (respectively $Q_{w}\cap\Pi^{\prime}=\emptyset$). A singular quadric threefold cannot contain a disjoint pair of planes, so the base loci of $f$ and $f^{\prime}$ disagree, as needed. ∎
@@ 5.3. Lattice theory and the movable cone
As in [flops, Proposition 7.2], the planes $\mathbb{P}(V)$ and $\mathbb{P}(V^{\vee})$ in $F$ yield walls $(g\pm 2\lambda)^{\perp}$ in $\mathrm{Mov}(F)$ on the boundary of $\mathrm{Nef}(F)$. The plane $P^{*}\subset F$ dual to $P$ yields a wall $(g-2p)^{\perp}$ also on the boundary of $\mathrm{Nef}(F)$. Further, the variety of lines incident to $P$ comprise a conic bundle over a K3 surface, and the wall $p^{\perp}$ lies on the boundary of both $\mathrm{Nef}(F)$ and $\mathrm{Mov}(F)$. Hence,
@@
Direct computation shows that $g$ is the unique class of degree $6$ pairing positively with all of $g\pm 2\lambda$, $g-2p$, and $p$. ∎
@@
We claim that $G_{2}\cdot\mathrm{Nef}(F_{1}^{\vee})=\mathrm{Nef}(F_{1}^{\vee})$. Since $\iota^{\vee}$ is regular on $F_{1}^{\vee}$, and $(\iota^{\vee})^{*}(g+\lambda)$ (verified using Lemma 5.12), we know $g+\lambda\in\mathrm{Nef}(F_{1}^{\vee})$. By calculating $G_{2}(g+\lambda)=g+\lambda$, we verify that $G_{2}$ sends $\mathrm{Nef}(F_{1}^{\vee})$ to itself.

### v2

@@ 1. Introduction
    Fourier-Mukai partnership, i.e. equivalence of the Kuznetsov components $\mathcal{A}_{X}$ of the derived categories of the two cubics, as defined in [Kuznet].
Motivated by the famous rationality conjectures for cubic fourfolds (see [hassett], [Kuznet], [addington], [MR4292740]), Huybrechts conjectured that (FM) implies (BE) [Huybrechtscubicsbook, Conjecture 7.3.21]. A very general cubic fourfold, i.e. $X\not\in\cup\,\mathcal{C}_{d}$, has no non-trivial Fourier-Mukai partners by [Huy17]. This conjecture becomes interesting for $X\in\mathcal{C}_{d},$ with some evidence provided by Fan and Lai [FL24] who verified the conjecture for a very general cubic in $\mathcal{C}_{20}$. We provide further evidence for this conjecture by considering two special families of cubics in $\mathcal{C}_{12}$ (see Theorems 3.1 and 5.1). The first is the family $\mathcal{M}_{nonsyz}$, parametrizing cubics containing a non-syzygetic pair of cubic scrolls, first studied in [BFMQ24]. The second is a component of $\mathcal{C}_{8}\cap\mathcal{C}_{12}$, first studied in [BolognesiRusso] and [yang2021lattice] and more recently in [C8C12]. On the other hand, it is well known that (BE) does not imply (FM), for example by considering two very general cubics in $\mathcal{C}_{14}$.
@@
It has been conjectured that (FM) implies (DF) [Huybrechtscubicsbook, Remark 7.3.28(iii)], for which we provide corroboratory evidence (see Theorems 3.1, 4.1, and 5.1). On the other hand, we prove that (DF) implies neither (FM) nor (BF), by considering the very general member of $\mathcal{C}_{32}$ (See Theorem 6.1), and $\mathcal{C}_{546}$ (see Theorem 7.1), respectively.
@@ 2.1. Special cubic fourfolds
have an associated K3 surface $S$, meaning $\widetilde{H}(\mathcal{A}_{X},\mathbb{Z})\cong\widetilde{H}(S,\mathbb{Z})$. All cubic fourfolds known to be rational belong to $\mathcal{C}_{d}$ for some $d$ satisfing ()(**)( ∗ ∗ ), leading to the famous conjecture that cubic fourfolds are rational if and only if they admit a labelling of disrciminant $d$ satisfying ()(**)( ∗ ∗ ). Kuznetsov conjectured in [kuzcubic] that $X$ is rational if and only if $\mathcal{A}_{X}\simeq D^{b}(S)$ for some K3 surface $S$, which was shown in [AddingtonTwoRatConjs] and [MR4292740] to coincide with the previous conjecture.
@@ 2.3. Fano varieties of lines
The variety $F$ is naturally equipped with the Plücker polarization $g\in\mathrm{NS}(F)$, with $g^{2}=6$ and divisibility $2$. Recall that $\mathcal{M}_{6}^{(2)}$ is the moduli space of hyperkähler fourfolds of K3[2] type equipped with a polarization $L$ with $L^{2}=6$ and divisibility $2$. There is a period map: $p\colon\mathcal{M}_{6}^{(2)}\rightarrow\mathcal{D}_{6}^{(2)},$ where $\mathcal{D}_{6}^{(2)}$ is the relevant period domain (see for instance [debarre2020hyperkahler, Appendix B]).
@@ 4.1. Lattice theory and the movable cone
Conversely, if $\varphi$ preserves the positive cone and has the prescribed action on $H,$ then one can extend $\varphi$ to an isometry $\tilde{\varphi}\in\mathrm{O}(H^{2}(F,\mathbb{Z}))$ (here we are using the fact that $q$ does not represent $-2$, so the positive cone and movable cone coincide). By construction, $\tilde{\varphi}$ preserves the Hodge structure, and hence $\varphi\in\mathrm{Im}(\pi\colon\mathrm{Mon}^{2}_{Hdg}(F)\rightarrow\mathrm{O}(\mathrm{NS}(F))).$ By Theorem 2.1, the conclusion follows. ∎
@@
As previously mentioned, $q$ does not represent $-2$, so $\mathrm{Mov}(F)=\overline{\mathrm{Pos}(F)}$. Let $\rho=ag+b\lambda$ where $106a^{2}+4ab-6b^{2}=-$ 10, so $\rho^{\perp}$ is a wall in $\mathrm{Mov}(F)$. If $|a|>1$, then $|13a+18b|<|a|$ or $|25a-18b|<|a|$, so exactly one of the classes $AB\cdot\rho$ or $(AB)^{-1}\cdot\rho$ has smaller first coordinate than $\rho$. Inductively, we see that $\rho=\pm(AB)^{n}\cdot(g+2\lambda)$ for a unique integer $n$. Hence $\mathrm{Bir}(F)$ acts freely and transitively on the walls of $\mathrm{Mov}(F)$. Further, $\mathrm{Bir}(F)$ acts transitively on the chambers of $\mathrm{Mov}(F)$ since $\mathrm{Mon}^{2}_{Hdg}\cong W_{Bir}\cong\mathrm{Bir}(F)$.
@@ 5.2. Geometry
  • since $[\ell]\not\in P^{*}$, $C_{\ell}\not\subset\Pi^{\prime}$, so the total transform of $C_{\ell}$ under the projection $Z_{T_{L}}\dashrightarrow X^{\prime}$ is a conic $D_{\ell}$.
Let $\ell^{\prime}$ be the line residual to $D_{\ell}$ in $X^{\prime}$. Moreover, let $C_{\ell^{\prime}}$ be the total transform of $\ell^{\prime}$ under the birational map $X^{\prime}\dashrightarrow Z_{T_{L}}$, and let $Q_{\ell}$ and $Q_{\ell^{\prime}}$ be the singular quadric threefolds associated to the conics $C_{\ell}$ and $C_{\ell^{\prime}}$ as in [BFMQ24, Lemma 2.22].
We first show that the formula $\pi([\ell])=[\ell^{\prime}]$ defines the same birational equivalence $\pi:F\dashrightarrow F^{\prime}$ as the composition of birational equivalences
[…]
defined in [BFMQ24, Proposition 3.6], where $\tau$ is the covering involution on $\widetilde{W}_{T_{L}}^{\perp}$. This amounts to checking that $Q_{\ell}=Q_{\ell^{\prime}}$ and that $C_{\ell}$ and $C_{\ell^{\prime}}$ span planes in opposite rulings of this quadric surface. Indeed, conics in the same ruling of a singular quadric threefold, or conics in two distinct singular quadric threefolds parametrized by $\widetilde{W}_{T_{L}}^{\perp}$ meet in at most one point, whereas $C_{\ell}$ and $C_{\ell^{\prime}}$ meet in two points, corresponding to the intersection of $D_{\ell}$ and $\ell^{\prime}$.
[…]
@@ 5.3. Lattice theory and the movable cone
As in [flops, Proposition 7.2], the planes $\mathbb{P}(V)$ and $\mathbb{P}(V^{\vee})$ in $F$ yield walls $(g\pm 2\lambda)^{\perp}$ in $\mathrm{Mov}(F)$ on the boundary of $\mathrm{Nef}(F)$. The plane $P^{*}\subset F$ dual to $P$ yields a wall $(g-2p)^{\perp}$ also on the boundary of $\mathrm{Nef}(F)$. Further, the variety of lines incident to $P$ composes a conic bundle over a K3 surface, and the wall $p^{\perp}$ lies on the boundary of both $\mathrm{Nef}(F)$ and $\mathrm{Mov}(F)$. Hence,
@@
Direct computation shows that $g$ is the unique class of degree $6$ in this cone pairing positively with all of $g\pm 2\lambda$, $g-2p$, and $p$. ∎
@@
We claim that $G_{2}\cdot\mathrm{Nef}(F_{1}^{\vee})=\mathrm{Nef}(F_{1}^{\vee})$. Since $\iota^{\vee}$ is regular on $F_{1}^{\vee}$, and $(\iota^{\vee})^{*}(g+\lambda)=g+\lambda$ (verified using Lemma 5.12), we know $g+\lambda\in\mathrm{Nef}(F_{1}^{\vee})$. By calculating $G_{2}(g+\lambda)=g+\lambda$, we verify that $G_{2}$ sends $\mathrm{Nef}(F_{1}^{\vee})$ to itself.